write an equation of the line containing the given point and parallel to the given line.Express your answer in the form y=mx+b. (3,7);x+2y=5
Answer by RAY100(1637)

x +2y = 5
.
2y = -x +5
.
y = (-1/2)x +(5/2),,,,,,,original line,,,m=-1/2, b=5/2,,
.
new parallel line has same slope
.
y=mx+b,,,,y= (-1/2)x +b
,
subst (3,7),,,,,7= (-1/2)(3) +b,,,,b=8.5
.
or new eqn is y=(-1/2)x +8.5
.
check,,,(3,7),,,7 = -1/2*3 +8.5 = 7 ,,,,,ok